OpenStudy (anonymous):

I have not been in school for 31 years and I am taking my first algebra course ever I do not understand it all I need help understanding algebra please

Directrix (directrix):

It all begins with your first problem. Ask.

OpenStudy (anonymous):

what does 5y(8+4) -8 mean

OpenStudy (anonymous):

first off 8 + 4 = 12 so it is the same as \[5y(12)-8\]

OpenStudy (anonymous):

when the variables are next to numbers with no symbol between them it measn multiplication, so this is the same as \[5\times y\times 12-8=5\times 12\times y-8=60y-8\]

OpenStudy (anonymous):

ok so would the answer be 52y

Directrix (directrix):

60 y - 8 =? 60y and 8 are not like terms and their difference cannot be further simplified. So, the answer is 60y -8. y is a variable and for all we know might be equal to, say 7. Then 60y - 8 = 60(7) -8 = 412. But, then y might be equal to 3 in which case 60y - 8 would be 60(3) - 8 = 180 - 8 = 172. Algebra can be thought of as arithmetic with variables or as some would say, "the science of finding x."